OpenStudy (anonymous):

what is 9 divided by 1/4?

OpenStudy (anonymous):

36. Invert the denominator then multiply.

OpenStudy (anonymous):

In general, to divide something by a/b, multiply it by b/a.

OpenStudy (anonymous):

jus multiply 9 nd 4 u will get the answer which is 36

OpenStudy (anonymous):

So what's 8 divided by 2/5?

OpenStudy (anonymous):

\[8 \div \frac{2}{5} \]\[= \frac{8}{\frac{2}{5}}\]\[=\frac{8}{\frac{2}{5}}\times 1\]\[=\frac{8}{\frac{2}{5}} \times \frac{5}{5}\]\[=\frac{8\times 5}{\frac{2}{5}\times 5}\]\[=\frac{40}{2} \]\[= 20\]

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

When dividing I prefer to write everything in fractional form, then simplify complex fractions by multiplying by various forms of 1 (as I did above with 1=5/5)

OpenStudy (anonymous):

I'm an old guy, bit lazy....

OpenStudy (anonymous):

estudier its 1/20

OpenStudy (anonymous):

Aargh!

OpenStudy (anonymous):

There's also the often taught method of.. \[a \div \frac{b}{c} = a \times \frac{c}{b}\] Which you can derive using my method.

OpenStudy (anonymous):

estudier its 20 i was wrong.........:(

OpenStudy (anonymous):

Which is what I posted to start with:-)

OpenStudy (anonymous):

Lol, oh. Sorry estudier.. I thought you were the OP. I take it you don't need help here ;)

OpenStudy (anonymous):

@ soniya show me your working...

OpenStudy (anonymous):

Note that soniya is also not the original poster.

OpenStudy (anonymous):

No matter, all wrong answers welcome (sort of).

OpenStudy (anonymous):

@soniya just to convince me, 15 divided by 3/4?

OpenStudy (anonymous):

actually iam new to thi site so dont know to post da eqution..iam trying to estudier

OpenStudy (anonymous):

no prob, just type it for now, you can play with the editor after...

OpenStudy (anonymous):

20 is da answer estudier

OpenStudy (anonymous):

am i correct

OpenStudy (anonymous):

OK, you seem to have got that, so 3/4 divided by 2/5?

OpenStudy (anonymous):

15 by 8 is da answer

OpenStudy (anonymous):

See, easy when you know how, isn't it?

OpenStudy (anonymous):

yes......